OpenStudy (anonymous):

Rebekah purchased a sweater that was reduced by25% (or 1/4) of its original price. She also used a store coupon for . When she made the purchase, she paid the cashier$19 . What was the original price of the sweater

OpenStudy (karatechopper):

How much was the store coupon?

OpenStudy (karatechopper):

Hello???

OpenStudy (karatechopper):

Vivi?

OpenStudy (anonymous):

yes i am here

OpenStudy (karatechopper):

ok! Can i ask how much the store coupon was?

OpenStudy (anonymous):

$5

OpenStudy (karatechopper):

ok!

OpenStudy (anonymous):

SO I WILL I BE GETTING THE ASWER I AM SO NOT GOOD IN MATH

OpenStudy (karatechopper):

You don't want an explanation on how to do the problem?

OpenStudy (anonymous):

YEAH U CAN GIVE EXPLANATION SO I CAN SEE HOW THE U GOT ANSWER

OpenStudy (karatechopper):

....ok.

OpenStudy (karatechopper):

So for this problem we will work backwards! We will start with $19 because that is what she paid. Now lets break the problem down. -She also used a store coupon for $5 So in this case we will add $5 to the 19. so 19+5=24 Now that we have the discounted price..we will take 1/4 of that to find out how much money was discounted. So... 24/4=6 Now we will add 6 to 24 for the original price. 24+6=30. Therefore, the original price of the Sweater is $30.

OpenStudy (karatechopper):

@vivi305